completely factor the expression y^2+12y+35
a. (y-7)(y-5)
b. (y+7)(y-5)
c. (y+7)(y+5)
d. prime

Answer by Edwin McCravy(20086)

FOIL 'em all out and see:

a. (y-7)(y-5) = y²-5y-7y+35 = y²-12y+35  No, the sign of -12y is wrong
b. (y+7)(y-5) = y²-5y+7y+35 = y²+2y+35  No, the middle term is wrong
c. (y+7)(y+5) = y²+5y+7y+35 = y²+12y+35  Yes, that's the same as the original.
d. prime.  No for c is the correct factorization.

Answer: c

Edwin